OpenStudy (anonymous):

A man weighs 750N on the surface of the earth. What would be his weight when standing on the moon? The masses of the and the are respectively, 5.98x10^24kg and 7.36x10^22kg. Their radii are respectively 6.37x10^3km and 1.74x10^3km

OpenStudy (anonymous):

you should post this question in physics group

OpenStudy (karatechopper):

yea i think its a good idea for physics place plz

OpenStudy (farmdawgnation):

successboye79, these guys are right. This does belong in the Physics group. I understand that there are more people here - but you still need to try and keep the question topic focused. Feel free to post in physics then post a link to your question in the chat rooms to get more help on it. Also, do not continue reposting the same question over-and-over. It's considered spamming, and is an offense that you can be suspended for. Best of luck finding answers to your questions. Feel free to holler at me if you have any questions about this.

OpenStudy (karatechopper):

holler,hehe

OpenStudy (anonymous):

thanks for that

OpenStudy (anonymous):

good for that

OpenStudy (anonymous):

i need solution to the question pls

OpenStudy (karatechopper):

post in physics then we can help u plz

OpenStudy (anonymous):

you will have to find gravity on the moon

OpenStudy (anonymous):

\[G m/r^2\] mass of the moon=7.36 * 10^22 kg radius of moon = 1 737.4 *10^3 meter

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{\left(6.67300\times 10^{-11}* 7.36\times 10{}^{\wedge}22\right)}{\left(1737.4*10^3\right)^2}\]

OpenStudy (farmdawgnation):

Thanks for helping him out imranmeah91. successboye79: I got this much attention for your question just by asking some people in the chat to help. So, you should be able to post in Physics and ask in Math chat for help if no one seems to be around. So, do that from now on please. :)

OpenStudy (anonymous):

solved out to be g=1.62704

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so we divide by earth's gravity and multiply by moon's gravity (750/9.8)(1.62704)=124.518

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

1/6 as much on the earth